what is the method to find kind of hybridisation in a compound in easy way 

Grade:12th Pass

2 Answers

Swapnil Saxena
12 years ago

There are certain tricks to effectively find the hybridization in a compound

1.No. of Hybridized orbitals = ( (No. of Valence Electrons in the central atom + No. of monovalet atoms surrounding the central alom (like H,F,Cl) - Charge of the species) )/2

Eg. NH4+  

---------------------------------------------------------------------------------------------------

No. of Valence Electrons in the central atom = 5

No. of monovalet atoms surrounding the central alom = 4

Charge of the species = +1

No. of hybridized orbital = ((5+4)-1)/2 = 4 (SP3)

This method works effectively for all compouns except the ones having single electrons loke NO2

------------------------------------------------------------------------------------------------------------------------------------

For this Purpose an another commonly used trick is

No. of Hybridized orbitals = No. of Atoms (surrounding the central atom) + No. of sigma bonds (excluding the dative bond) + No of lone pairs

Eg. NO2

No. of Atoms = 2 (two O)

No. of Sigma Bonds= 1

No of lone pairs = 0

No. of Hybridized orbitals = 3 (sp2)

or

No. of Hybridized orbitals = No. of sigma bond + No. of dative bond + No of lone pairs + No of single electrons


 






suchita undare
12 years ago

no. of lone pairs+no. of bond pairs

 e.g BeCl2 no.of lone pairs=0 no of bond pairs=2 therefore 2+0=2

therefore sp hybridisation